OpenStudy (anonymous):

solve for x: 9^(2x+1) = 27^(x-2)

OpenStudy (anonymous):

\[9^{2x+1} = 27^{x-2}\]

ganeshie8 (ganeshie8):

hint : powers of 3

OpenStudy (anonymous):

yes so what next?

OpenStudy (john_es):

\[9^{2x+1}=27^{x-2}\Rightarrow (3^2)^{2x+1}=(3^3)^{x-2}\Rightarrow 3^{4x+2}=3^{3x-6}\] \[\Rightarrow 4x+2=3x-6\] And solve.

OpenStudy (anonymous):

ok thanks mate